# Hybrid Video Diffusion Models with 2D Triplane and 3D Wavelet Representation

### v1

@@ Abstract
Generating high-quality videos that synthesize desired realistic content is a challenging task due to their intricate high-dimensionality and complexity of videos. Several recent diffusion-based methods have shown comparable performance by compressing videos to a lower-dimensional latent space, using traditional video autoencoder architecture. However, such method that employ standard frame-wise 2D and 3D convolution fail to fully exploit the spatio-temporal nature of videos. To address this issue, we propose a novel hybrid video diffusion model, called HVDM, which can capture spatio-temporal dependencies more effectively. The HVDM is trained by a hybrid video autoencoder which extracts a disentangled representation of the video including: (i) a global context information captured by a 2D projected latent (ii) a local volume information captured by 3D convolutions with wavelet decomposition (iii) a frequency information for improving the video reconstruction. Based on this disentangled representation, our hybrid autoencoder provide a more comprehensive video latent enriching the generated videos with fine structures and details. Experiments on video generation benchamarks (UCF101, SkyTimelapse, and TaiChi) demonstrate that the proposed approach achieves state-of-the-art video generation quality, showing a wide range of video applications (e.g., long video generation, image-to-video, and video dynamics control).
${}^{\dagger}$orresponding author
@@ 1 Introduction
Figure 1: Overview of our hybrid video autoencoder, called HVDM. The HVDM combines a 2D triplane and 3D volume representation for video encoding. The 2D triplane representation provide global context and 3D volume representation provide local volume information of video. The spatio-temporal cross-attention module incorporates these distinctive features to organizes fine-grained video representation.
@@
By leveraging the recent advancements of image-based diffusion models [15, 14], a variety of video diffusion methods have emerged [17, 28, 25]. These methods show considerable promise in modeling video distributions and synthesizing visually plausible results. However, they often face efficiency challenge due to operating in pixel space. To address this challenge, several recent latent video diffusion models have been developed, training within a compressed latent space to reduce computational demands [51, 13, 10, 51, 2]. While these latent diffusion models exhibits improved scaling properties with reduced complexity, the potential of novel autoencoder designs remains largely unexplored. There exist conventional approaches, employing standard 2D or 3D CNN-based autoencoders for video generation [17, 31, 13, 19]. Yet, these CNN-based architectures struggle with effectively managing the overall spatio-temporal structure of videos. While transformers might be considered as a solution, their direct application to video representation learning can lead to significant computational time and memory inefficiency [21, 44].
@@
Unlike the aforementioned approaches, PVDM [48] offers a novel perspective in video generation. PVDM simplifies video complexity by adopting a triplane representation, where video data is factorized into 3D projected latents across different spatio-temporal directions in a latent space. This method has demonstrated improved generation results over previous works such as StyleGAN-based approaches [32] and other diffusion models [17], particularly in effectively representing intricate video data. Despite its progress, they encounter inherent limitations due to the reduction of higher-dimensional video data into lower-dimensional spaces. The 2D projected latents, extracted by transformer, struggle to fully recover the 3D structure of video volume, lacking crucial volume information. Furthermore, while transformers have the potential to achieve significant improvements when trained with larger datasets and increased model capacity, nonetheless, it seems unnecessary to overlook the established understanding of video properties, such as spatial and temporal redundancy [11]. Rather than reinventing the wheel, leveraging these known characteristics could enhance efficiency and effectiveness in video generation tasks.
@@
To effectively utilize both global context and local volume information in video, we propose a novel video autoencoder designed to comprehensively capture the spatio-temporal dependencies of video. Our video autoencoder combines the transformer and 3D CNN architectures to obtain dense video representation. We start from a well-known observation that transformers excel at capturing long-range dependencies, and we employ them to acquire the global context of the video. To compensate the limited inductive bias of transformers, we leverage 3D CNNs to provide local volume details, drawing from the recognized advantage of 3D CNNs in capturing local spatial information. More specifically, our proposed autoencoder combines the stengths of both architectures by incorporating global 2D projected and local 3D volume latent representations. These two types of latent representations mutually enhance each other through cross-attention module that operates along both the temporal and spatial axes.
@@
Taking this approach further, we incorporate wavelet transformation for more nuanced video volume encoding. Using 3D discrete wavelet transform that provides a rich representation for video in the spatio-spectral domain, we can decompose complex video into different frequency components at multiple levels of detail. We observe that incorporating wavelet-based features enhances model convergence, compared to using the raw video data. Additionally, we enhanced video realism by training this video autoencoder with both reconstruction and frequency matching loss. Our frequency matching loss ensures the preservation of both coarse and fine frequency components and leads to improved visual quality during video reconstruction process.
@@ 2 Related Work
Tri-plane representations
The effectiveness of tri-plane representation has already found in 3D-aware generation works [29, 8, 7, 3, 43, 12]. These tri-plane representation methods have been shown to offer an effective compression benefits by projecting the 3D volume as 2D-shaped latent features. By taking advantage of these tri-plane representation, PVDM  [48] achieves comparable results in the video generation task, encoding the cubic structure of video pixels. Inspired by this, we present an improved video encoding method that exploit the 2D projected latent with wavelet-based 3D volume latent.
@@ 3 Preliminaries
3.1 2D Triplane for Video Diffusion Model
Given a video sample $\mathbf{x}\in\mathbb{R}^{3\times S\times H\times W}$ with time $S$, height $H$, and width $W$, the tri-plane autoencoder encodes it as three 2D projected latent $[\mathbf{z}_{\mathrm{proj}}^{\mathrm{hw}},\mathbf{z}_{\mathrm{proj}}^{\mathrm{ws}},\mathbf{z}_{\mathrm{proj}}^{\mathrm{hs}}]$ where $\mathbf{z}_{\mathrm{proj}}^{\mathrm{hw}}\in\mathbb{R}^{C\times H^{\prime}\times W^{\prime}}$, $\mathbf{z}_{\mathrm{proj}}^{\mathrm{ws}}\in\mathbb{R}^{C\times S\times W^{\prime}}$, $\mathbf{z}_{\mathrm{proj}}^{\mathrm{hs}}\in\mathbb{R}^{C\times S\times H^{\prime}}$. We denote $C$ as number of channels, $H^{\prime}=H/d$, $W^{\prime}=W/d$ for downsampling factors $d>1$. Unlike conventional video autoencoder which depends on frame-wise 2D or 3D convolution networks [17, 31, 13, 19], the PVDM [48] effectively decompose the 3D data using video transformer [4] for projections $[\mathbf{z}_{\mathrm{proj}}^{\mathrm{hw}}$, $\mathbf{z}_{\mathrm{proj}}^{\mathrm{ws}}$, $\mathbf{z}_{\mathrm{proj}}^{\mathrm{hs}}]$. The utilization of these latent vectors $[\mathbf{z}_{\mathrm{proj}}^{\mathrm{hw}},\mathbf{z}_{\mathrm{proj}}^{\mathrm{ws}},\mathbf{z}_{\mathrm{proj}}^{\mathrm{hs}}]$ offers several advantages in terms of achieving both computation and memory efficiency. Furthermore, their image-like structure enables the efficient video generation from the latent space reducing the dimensionality of the data.
3.2 3D Wavelet Transform
3D Wavelet transform is a classical algorithm widely used in medical imaging and video compression to separate different frequency components within the data. Just as the 2D wavelet transform, given $\mathbf{x}\in\mathbb{R}^{3\times S\times H\times W}$, the 3D wavelet transform operates along $s$, $h$, and $w$ direction decomposing the 3D volume into different subbands. These subbands are factorized into low and high subbands where the low subband express the slowly changing part of the data and the high subband contain fine details capturing sharp edges and rapid transitions within the data.
@@ 4.1 Designing Hybrid Video Autoencoder
2D triplane representation. The 2D projected latent which is originated from PVDM [48] is extracted by 3D-to-2D projections mapping at each spatio-temporal direction to encode 3D video pixels. As shown in Sec 3.1, the 2D image-like latent $[\mathbf{z}_{\mathrm{proj}}^{\tt hw},\mathbf{z}_{\mathrm{proj}}^{\tt ws},\mathbf{z}_{\mathrm{proj}}^{\tt hs}]$ can be encoded with a combination of a video encoder [4] and a small transformer [40]. Formally, the projected latent is formulated as follows:
@@
where $\mathcal{T}_{\tt video}^{\mathrm{shw}}$ denotes a video-to-3D-latent encoder such as video transformer [4] and $\mathcal{T}_{\mathrm{proj}}=[\mathcal{T}_{\mathrm{proj}}^{\mathrm{s}},\mathcal{T}_{\mathrm{proj}}^{\mathrm{h}},\mathcal{T}_{\mathrm{proj}}^{\mathrm{w}}]$ denotes a small transformer [40] for projections. While projected latent provide a global information of video with a computational efficiency, this transformer architecture projecting a triplane leads to lossy compression due to the lack of local volume information. To overcome this limitation, we transformed the 2D projected latent into a 3D volume and complemented it with the 3D volume representation. Converting a 2D projected latent into a 3D volume $\mathbf{z}_{\tt proj}$ is formulated as follows:
@@
3D wavelet representation. 3D CNN based volume encoder $\mathcal{E}_{\mathrm{low}},\mathcal{E}_{\mathrm{high}}$ extract local spatio-temporal features in videos. These volumetric local features are exploited with the 2D projected latent $\mathbf{z}_{\mathrm{proj}}$ to reorganize fine-grained features. Instead of utilizing the unprocessed RGB video frame directly in the volume encoder, we employ the decomposed wavelet features as input. These decomposed wavelet features extracted from the 3D wavelet transform $\mathcal{W}$ provide coarse and detailed volume coefficient suited for video encoding. Extracting 3D wavelet latent can be formulated as follows:
@@
Fusing 2D triplane and 3D wavelet representations. As shown in Fig. 1, these 2D projected and 3D volume latent are integrated by cross-attention operation across the temporal and spatial axis. In this way, we can reinforce the volume-related inductive bias and rearrange fine-grained video latent for video generation. Formally, this cross-attention process can be formulated as follows:
@@
Figure 3: The framework of our proposed HVDM. Our HVDM is adaptable for diverse video generation tasks depending on the type of condition latent $z_{c}$. During training process, the condition latent $z_{c}$ is extracted from the video and jointly trained with the noised hybrid video latent $z_{t}$. During the sampling process, these condition latent are supported to enable various video generation tasks, such as long video generation, image-to-video, and video dynamics control.
@@ 4.2 Loss function
Reconstruction loss. The reconstruction loss $\mathcal{L}_{rec}$ denotes the similarity between the original input video and output video reconstructed by a model. The reconstruction loss term enforce the model to learn meaningful features in the encoding phase so that it can accurately reconstruct the input data in the decoding phase. We employ mean absolute error for reconstruction:
@@
Perceptual loss. The perceptual loss measures the difference between the feature map of input and a generated output in terms of perceptual similarity. Typically, this loss entails the extraction of feature maps from a pre-trained neural network. In contrast to reconstruction loss, which focuses solely on the pixel-level differences between videos, perceptual loss takes into account high level features such as textures and structures that humans perceive.
@@
Frequency matching loss. Conventional methods typically rely on pixel-level losses in the spatial-temporal space. Since these loss functions closely match the original videos in pixel domain, the generated video’s quality may not be sufficient potentially failing to capture the frequency information. Our frequency matching loss is designed to facilitate the model can learn frequency features with less optimization difficulty. This loss not only helps the model to recover missing frequency components but also enhances the overall reconstruction quality. We define the frequency matching loss as a reconstruction term in the wavelet domain:
@@
Overall loss. Taking into account all the losses, the overall loss is formulated as follows:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\mathrm{rec}}+\lambda_{2}\mathcal{L}_{\mathrm{lpips}}+\lambda_{3}\mathcal{L}_{\mathrm{freq}}$ (9)
where we choose $\lambda_{1}=\lambda_{2}=1$ and $\lambda_{3}=0.5$ in our experiments.

### v4

# https://hxngiee.github.io/HVDM/

@@ Abstract
Generating high-quality videos that synthesize desired realistic content is a challenging task due to their intricate high dimensionality and complexity. Several recent diffusion-based methods have shown comparable performance by compressing videos to a lower-dimensional latent space, using traditional video autoencoder architecture. However, such method that employ standard frame-wise 2D or 3D convolution fail to fully exploit the spatio-temporal nature of videos. To address this issue, we propose a novel hybrid video diffusion model, called HVDM, which can capture spatio-temporal dependencies more effectively. HVDM is trained by a hybrid video autoencoder which extracts a disentangled representation of the video including: (i) a global context information captured by a 2D projected latent, (ii) a local volume information captured by 3D convolutions with wavelet decomposition, and (iii) a frequency information for improving the video reconstruction. Based on this disentangled representation, our hybrid autoencoder provide a more comprehensive video latent enriching the generated videos with fine structures and details. Experiments on standard video generation benchmarks such as UCF101, SkyTimelapse, and TaiChi demonstrate that the proposed approach achieves state-of-the-art video generation quality, showing a wide range of video applications (e.g., long video generation, image-to-video, and video dynamics control). The source code and pre-trained models will be publicly available once the paper is accepted.
footnotetext: Corresponding author
@@ 1 Introduction
By leveraging the recent advancements of image-based diffusion models [16, 15], a variety of video diffusion methods have emerged [18, 30, 27]. These methods show considerable promise in modeling video distributions and synthesizing visually plausible results. However, they often face efficiency challenge due to operating in pixel space. To address this challenge, several recent latent video diffusion models have been developed, training within a compressed latent space to reduce computational demands [56, 14, 12, 56, 2]. While these latent diffusion models exhibit improved scaling properties with reduced complexity, the potential of novel autoencoder designs remains largely unexplored. There exist conventional approaches, employing standard 2D or 3D CNN-based autoencoders for video generation [18, 34, 14, 20]. Yet, these CNN-based architectures struggle with effectively managing the overall spatio-temporal structure of videos. While transformers might be considered as a solution, their direct application to video representation learning can lead to significant computational time and memory inefficiency [22, 49].
@@
Unlike the aforementioned approaches, PVDM [53] offers a novel perspective in video generation. PVDM simplifies video complexity by adopting a triplane representation, where video data is factorized into 2D projected latents across different spatio-temporal directions in a latent space. This method has demonstrated improved generation results over previous works such as StyleGAN-based approaches [35] and other diffusion models [18], particularly in effectively representing intricate video data. Despite its progress, they encounter inherent limitations due to the reduction of higher-dimensional video data into lower-dimensional spaces. More specifically, the 2D projected latents struggle to fully recover the 3D structure of video volume, lacking crucial volume information.
@@
Motivated by these observations, we propose a novel hybrid video diffusion model designed to comprehensively capture the spatio-temporal dependencies of video, called HVDM. Rather than reinventing the wheel, we leverage these known characteristics to enhance effectiveness in video generation tasks. Our video autoencoder combines transformer for 3D-to-2D projections and the 3D CNNs to effectively utilize both global context and local volume information in video. We first start from a well-known fact that transformers excel at capturing long-range dependencies, and we employ them to acquire the global context of the video. To compensate the limited inductive bias of transformers, we leverage 3D CNNs to provide local volume details, drawing from the recognized advantage of 3D CNNs in capturing short-range spatio-temporal information. More specifically, our proposed autoencoder builds the stengths of both architectures by incorporating global 2D projected and local 3D volume latent representations. These two types of latent representations mutually enhance each other through cross-attention module that operates along both the temporal and spatial axes.
Figure 1: Overview of our hybrid video autoencoder in HVDM that combines a 2D triplane and 3D volume representation for video encoding. The 2D triplane representation provide global context and 3D volume representation provide local volume information of video. The spatio-temporal cross-attention module incorporates these distinctive features to organizes fine-grained video representation.
@@
Taking this approach further, we pose the question of whether it is possible to exploit frequency information for boosting the video reconstruction process. We find that employing 3D discrete wavelet transform could potentially resolve the problem of limited receptive fields in 3D CNNs. Specifically, considering that the 3D discrete wavelet transform reduces the video’s size by half in the frequency domain, it allows us to expand the receptive field without losing information. Additionally, using 3D discrete wavelet transform decompose complex video into different frequency components at multiple levels of details, providing rich video representation. Finally, by training the video autoencoder with both reconstruction and frequency matching loss in the pixel and spectral spaces, frequency matching loss improved visual quality during video reconstruction process. As a result, we observe that the integration of wavelet-based features provides benefit compared to the use of raw video data by enabling more nuanced video encoding.
@@ 2 Related Work
2.0.3 Triplane representations.
The effectiveness of triplane representation has already found in 3D-aware generation works [31, 9, 8, 3, 48, 13]. These triplane representation methods have been shown to offer an effective compression benefits by projecting the 3D volume as 2D-shaped latent features. By taking advantage of these triplane representation, PVDM  [53] achieves comparable results in the video generation task, encoding the cubic structure of video pixels. Inspired by this, we present an improved video encoding method that exploit the 2D projected latent with wavelet-based 3D volume latent.
@@ 3 Preliminaries
3.1 2D Triplane Representation for Video Diffusion Model
Given a video sample $\mathbf{x}\in\mathbb{R}^{3\times S\times H\times W}$ with time $S$, height $H$, and width $W$, the triplane autoencoder encodes it as three 2D projected latent $[\mathbf{z}_{\mathrm{proj}}^{\mathrm{hw}},\mathbf{z}_{\mathrm{proj}}^{\mathrm{ws}},\mathbf{z}_{\mathrm{proj}}^{\mathrm{hs}}]$ where $\mathbf{z}_{\mathrm{proj}}^{\mathrm{hw}}\in\mathbb{R}^{C\times H^{\prime}\times W^{\prime}}$, $\mathbf{z}_{\mathrm{proj}}^{\mathrm{ws}}\in\mathbb{R}^{C\times S\times W^{\prime}}$, $\mathbf{z}_{\mathrm{proj}}^{\mathrm{hs}}\in\mathbb{R}^{C\times S\times H^{\prime}}$. We denote $C$ as number of channels, $H^{\prime}=H/d$, $W^{\prime}=W/d$ for downsampling factors $d>1$. Unlike conventional video autoencoder which depends on frame-wise 2D or 3D convolution networks [18, 34, 14, 20], the PVDM [53] effectively decompose the 3D data using video transformer [5] for projections $[\mathbf{z}_{\mathrm{proj}}^{\mathrm{hw}}$, $\mathbf{z}_{\mathrm{proj}}^{\mathrm{ws}}$, $\mathbf{z}_{\mathrm{proj}}^{\mathrm{hs}}]$. The utilization of these latent vectors $[\mathbf{z}_{\mathrm{proj}}^{\mathrm{hw}},\mathbf{z}_{\mathrm{proj}}^{\mathrm{ws}},\mathbf{z}_{\mathrm{proj}}^{\mathrm{hs}}]$ offers several advantages in terms of achieving both computation and memory efficiency. Furthermore, their image-like structure enables the efficient video generation from the latent space reducing the dimensionality of the data.
3.2 3D Wavelet Representation
3D Wavelet transform is a classical algorithm widely used in medical imaging and video compression to separate different frequency components within the data [47, 46, 32]. Just as the 2D wavelet transform, given $\mathbf{x}\in\mathbb{R}^{3\times S\times H\times W}$, the 3D wavelet transform operates along $s$, $h$, and $w$ direction decomposing the 3D volume into different subbands. These subbands are factorized into low and high subbands where the low subband express the slowly changing part of the data and the high subband contain fine details capturing sharp edges and rapid transitions within the data.
@@ 4.1 Designing Hybrid Video Autoencoder
4.1.1 2D triplane representation.
The 2D projected latent which is originated from PVDM [53] is extracted by 3D-to-2D projections mapping at each spatio-temporal direction to encode 3D video pixels. As shown in Sec 3.1, the 2D image-like latent $[\mathbf{z}_{\mathrm{proj}}^{\tt hw},\mathbf{z}_{\mathrm{proj}}^{\tt ws},\mathbf{z}_{\mathrm{proj}}^{\tt hs}]$ can be encoded with a combination of a video encoder [5] and a small transformer [43]. Formally, the projected latent is formulated as follows:
@@
where $\mathcal{T}_{\tt video}^{\mathrm{shw}}$ denotes a video-to-3D-latent encoder such as video transformer [5] and $\mathcal{T}_{\mathrm{proj}}=[\mathcal{T}_{\mathrm{proj}}^{\mathrm{s}},\mathcal{T}_{\mathrm{proj}}^{\mathrm{h}},\mathcal{T}_{\mathrm{proj}}^{\mathrm{w}}]$ denotes a small transformer [43] for projections. While projected latent provide a global information of video with a computational efficiency, this transformer architecture projecting a triplane leads to lossy compression due to the lack of local volume information. Moreover, the weak inductive bias of the transformer leads to a slower convergence speed for the model. To overcome these limitations, we transformed the 2D projected latent into a 3D volume and complemented it with the 3D volume representation. Converting a 2D projected latent into a 3D volume $\mathbf{z}_{\tt proj}$ is formulated as follows:
@@
4.1.2 3D wavelet representation.
3D CNN based volume encoder $\mathcal{E}_{\mathrm{low}},\mathcal{E}_{\mathrm{high}}$ extract local spatio-temporal features in videos. These volumetric local features are exploited with the 2D projected latent $\mathbf{z}_{\mathrm{proj}}$ to reorganize fine-grained features. Instead of utilizing the unprocessed RGB video frame directly in the volume encoder, we employ the decomposed wavelet features as input. These decomposed wavelet features extracted from the 3D wavelet transform $\mathcal{W}$ provide coarse and detailed volume coefficient suited for video encoding. Extracting 3D wavelet latent can be formulated as follows:
@@
4.1.3 Fusing 2D triplane and 3D wavelet representations.
As shown in Fig. 1, these 2D projected and 3D volume latent are integrated by cross-attention operation across the temporal and spatial axis. In this way, we can reinforce the volume-related inductive bias and rearrange fine-grained video latent for video generation. Formally, this cross-attention process can be formulated as follows:
@@ 4.2 Loss function
4.2.1 Reconstruction loss.
The reconstruction loss $\mathcal{L}_{rec}$ denotes the similarity between the original input video and output video reconstructed by a model. The reconstruction loss term enforce the model to learn meaningful features in the encoding phase so that it can accurately reconstruct the input data in the decoding phase. We employ mean absolute error for reconstruction:
@@
4.2.2 Perceptual loss.
The perceptual loss measures the difference between the feature map of input and a generated output in terms of perceptual similarity. Typically, this loss entails the extraction of feature maps from a pre-trained neural network. In contrast to reconstruction loss, which focuses solely on the pixel-level differences between videos, perceptual loss takes into account high level features such as textures and structures that humans perceive.
@@
4.2.3 Frequency matching loss.
Conventional methods typically rely on pixel-level losses in the spatial-temporal space. Since these loss functions closely match the original videos in pixel domain, the generated video’s quality may not be sufficient potentially failing to capture the frequency information. Our frequency matching loss is designed to facilitate the model can learn frequency features with less optimization difficulty. This loss not only helps the model to recover missing frequency components but also enhances the overall reconstruction quality. We define the frequency matching loss as a reconstruction term in the wavelet domain:
@@
4.2.4 Overall loss.
Taking into account all the losses, the overall loss is formulated as follows:
$\mathcal{L}=\lambda_{\mathrm{rec}}\mathcal{L}_{\mathrm{rec}}+\lambda_{\mathrm{lpips}}\mathcal{L}_{\mathrm{lpips}}+\lambda_{\mathrm{freq}}\mathcal{L}_{\mathrm{freq}},$ (9)
where $\lambda_{\mathrm{rec}}=\lambda_{\mathrm{lpips}}=1$ and $\lambda_{\mathrm{freq}}=0.5$ in our experiments.
@@
4.3 Diffusion Model-based Video Generator
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i})$ and $\beta$ represents a pre-defined schedule for the variance.
where $\epsilon_{\theta}(\cdot)$ is a trainable denoising diffusion model which recover a noisy sample $\mathbf{z}_{t}$. The denoising diffusion model is trained to predict the noise with the following simplified objective:
Figure 3: Applications for diverse video generator by our proposed HVDM. Our HVDM is adaptable for diverse video generation tasks depending on the type of condition latent $z_{c}$. During training process, the condition latent $z_{c}$ is extracted from the video and jointly trained with the noised hybrid video latent $z_{t}$. During the sampling process, these condition latent are supported to enable various video generation tasks, such as long video generation, image-to-video, and video dynamics control.
4.4 Applications for Diverse Video Generator
4.4.1 Long video generation.
To facilitate long video generation, our work follows the traditional frame prediction approach [44, 14, 53] where previous frame conditions are combined with future frames. Given two consecutive video clips $[\mathbf{x}^{\mathrm{prev}},\mathbf{x}^{\mathrm{next}}]$ of $\mathbf{x}^{\mathrm{prev}},\mathbf{x}^{\mathrm{next}}\in\mathbb{R}^{3\times S\times H\times W}$, we can train the model to generate initial clips and future clips in a unified manner. More specifically, we can generate an initial clip with a null latent and extend the video by composing the condition with previous frame latent as follow:
$\mathbf{z}_{c}^{\mathrm{long}}=\begin{cases}\mathcal{E}(\mathbf{x}^{\mathrm{prev}}),&~{}\text{if future clip},\\ 0,&~{}\text{otherwise},\end{cases}$ (14)
where $\mathcal{E}$ denotes our pre-trained hybrid encoder.
4.4.2 Image-to-video.
Image-to-video generation aims to synthesize a coherent video starting from an image. In general, videos are composed of a series of consecutive images, where the first frame serves as a representative depiction of the overall content within the video. To facilitate the process of image-to-video generation without introducing an additional image encoder, we repeatedly stack the first frame of the video on the temporal axis. Then, we use the latent $\mathcal{I}^{\mathrm{content}}$ extracted from our video encoder as the content condition. By leveraging the latent $\mathcal{I}^{\mathrm{content}}$ which capture the common content, we can effectively guide the model to generate coherent video sequence as follow:
4.4.3 Video dynamics control.
Empowering users to control motion intensity in video creation provides flexibility with an improved user experience. Unfortunately, it was difficult to obtain motion-related labels from existing video collections for training diffusion models. To enable the alteration of motion intensity, we devise a method that learns motion dynamics with inter-frame variance. Specifically, we define the level of motion alteration as the structural dissimilarity between the semantic latent of the start and end frames. Since the structural dissimilarity encompasses both semantic and motion variance, we employ it as guiding cues to control motion without explicit motion labeled video. Formally, our structural dissimilarity is defined as follow:
$\mathbf{z}_{c}^{\mathrm{dynamics}}=\|\mathbf{z}_{0}^{\tt T}-\mathbf{z}_{0}^{\tt 1}\|_{1}.$ (16)
$\mathbb{E}_{\mathbf{x}_{0},\epsilon,t}\Big{[}\lambda||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},\mathbf{z}_{c},t)||_{2}^{2}+(1-\lambda)||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},{0},t)||_{2}^{2}\Big{]},$ (17)
5 Experiments
5.1 Training Details
Datasets. We conduct the experiments on three benchmark datasets for the video generation: UCF-101 [38], SkyTimelapse [51], and TaiChi [33]. UCF-101 [38] contains 13,320 videos from 101 action classes, offering a diverse range of human actions. SkyTimelapse [51] is composed of 5,000 videos containing dynamic sky scenes. TaiChi [33] consists of full human bodies performing Tai-Chi actions.